# Structure-Aware Correspondence Learning for Relative Pose Estimation

### v1

@@ 4.1 Comparison with State-of-the-Art Methods
Our approach significantly outperforms traditional 2D feature-based methods, such as SuperGlue [33] and LoFTR [36], primarily due to their inability to reliably match keypoints under large pose differences and minimal overlap areas. Furthermore, our method achieves superior results compared to hypothesis-and-verification-based methods like RelPose [48] and RelPose++ [19]. These methods rely on global features while ignoring local structural cues, and their use of discrete sampling limits their ability to model the continuous pose space accurately.
@@ 4.3 Visualization
Qualitative Results. The qualitative results of LoFTR [36], DVMNet [50] and our method on the LineMOD dataset are shown in Figure 3. Specifically, we visualize the object pose depicted in the query image. The query object pose is determined as $\mathbf{R}_{q}=(\Delta\mathbf{R})^{-1}\mathbf{R}_{r}$, where $\mathbf{R}_{r}$ denotes the object pose in the reference image. The green and blue arrows represent the visualization of $\mathbf{R}_{q}$ calculated from the prediction and the ground truth, respectively. It can be observed that our method demonstrates more accurate performance compared to previous matching-based methods, especially in cases where there are larger viewpoint differences between the query and reference images.

### v2

@@ 4.1 Comparison with State-of-the-Art Methods
Our approach significantly outperforms traditional 2D feature-based methods, such as SuperGlue [sarlin2020superglue] and LoFTR [sun2021loftr], primarily due to their inability to reliably match keypoints under large pose differences and minimal overlap areas. Furthermore, our method achieves superior results compared to hypothesis-and-verification-based methods like RelPose [zhang2022relpose] and RelPose++ [lin2024relpose++]. These methods rely on global features while ignoring local structural cues, and their use of discrete sampling limits their ability to model the continuous pose space accurately.
@@ 4.3 Visualization
Qualitative Results. The qualitative results of LoFTR [sun2021loftr], DVMNet [zhao2024dvmnet] and our method on the LineMOD dataset are shown in Figure 3. Specifically, we visualize the object pose depicted in the query image. The query object pose is determined as $\mathbf{R}_{q}=(\Delta\mathbf{R})^{-1}\mathbf{R}_{r}$, where $\mathbf{R}_{r}$ denotes the object pose in the reference image. The green and blue arrows represent the visualization of $\mathbf{R}_{q}$ calculated from the prediction and the ground truth, respectively. It can be observed that our method demonstrates more accurate performance compared to previous matching-based methods, especially in cases where there are larger viewpoint differences between the query and reference images.